Refer to the accompanying table in answering the questions that follow: Aggregate Expenditures (Catlg+Xn+G), Billions 420 Real Domestic Output, Possible Levels of Employment, Millions 70 90 110 130 150 Billions 400 450 460 500 540 580 600 a. If full employment in this economy is 150 million, will there be an inflationary expenditure gap or a recessionary expenditure gap? (Click to select) What will be the consequence of this gap? (Click to select) By how much would aggregate expenditures in column 3 have to change at each level of GDP to eliminate the inflationary Aggregate expenditures would have to (Click to select) by$ billion What is the multiplier in this example? expenditure gap or the recessionary expenditure gap? b. Will there be an inflationary expenditure gap or a recessionary expenditure gap if the full-employment level of output is $400 billion? (Cick to select) By how much would aggregate expenditures in column 3 have to change at each level of GDP to eliminate the gap? Aggregate expenditures would have to (Click to select) by $billion. What is the multiplier in this example? ((click to select)
what is the multiplier in this example? C. Assuming that investment, net exports, and government expenditures do not change with changes in real GDP, what are the sizes the MPC, the MPS, and the multiplier? Instructions: Round your answers for MPC and MPS to 1decimal place. MPS = Multiplier-
Answer #1

A) Equilibrium level of GDP=500 Billion because aggregate expenditure= Real domestic output

Thus there will be recessionary gap of $100billion

Thus aggregate expenditure would have to increase by $20billion

Multiplier in this example 100/20=5

B) Full employment output=400billion

thus an inflationary Gap of $100billion

Aggregate expenditure has to decrease by $20 billion

again multiplier=100/20=5

c)MPC=40/50=0.8

MPS=1-MPC=1-0.8=0.2

Mulriplier=1/1-Mpc=1/1-0.8=1/0.2=5